OpenStudy (anonymous):

the complement of angle a is 38 degrees more than angle a find the measure of angle a please help

OpenStudy (ddcamp):

Compliment means the angles add up to 90°. If c is the compliment, then we have: a+c=90°

OpenStudy (ddcamp):

We're told that the complement is 38° more than the angle. Using that information, we have: a + (a+38) = 90 2a + 38 = 90

OpenStudy (anonymous):

do i subtract 38 on both sides?

OpenStudy (ddcamp):

Yes. Then divide by 2.

OpenStudy (anonymous):

so 26

OpenStudy (anonymous):

a=26

OpenStudy (ddcamp):

Yes

OpenStudy (anonymous):

thanks